Required information The JMH Company manufactures resistors with an ideal resistance value of 400 ohms. A sample of 7000 resistors were tested and found to have a mean resistance of 391.5 ohms with a standard deviation of 31.7 ohms. Assume a normal distribution. Given a tolerance of ±10 percent, what is the probability that a particular resistor will be unusable? Round to four decimal places. The probability that a particular resistor will be unusable is
Answer #1

here for tolerance -/+10% ; a resistor will be unusable if resistor have resistance below 360 or more then 440

hence P(resistor will be unusable) =P(X<360)+P(X>440)=1-P(360<X<440)

=1-P((360-391.5)/31.7<Z<(440-391.5)/31.7)=1-P(-0.99<Z<1.53)=1-(0.9370-0.1611)=0.2241